Four plans for repayment of $5000 in five years with interest at 8%
1) At end of each year, pay $1000 principal plus interest due.


2) Pay interest due at end of each year and principal at end of five years.


3) Pay in five equal end-of-year payments.


4) Pay principal and interest in one payment at end of five years
Answer #1

Plan A:

1st year payment=1000+5000*8%=1400

2nd year payment=1000+4000*8%=1320

3rd year payment=1000+3000*8%=1240

4th year payment=1000+2000*8%=1160

5th year payment=1000+1000*8%=1080

This plan has the least interest

Plan B:

1st year payment=5000*8%=400

2nd year payment=5000*8%=400

3rd year payment=5000*8%=400

4th year payment=5000*8%=400

5th year payment=5000+5000*8%=5400

Plan C:

Each year payment=5000*8%/(1-1/1.08^5)=1252.28

Plan D:

No payment in year 1,2,3,4

5th year payment=5000*1.08^5=7346.64

This plan has the most interest component